# Effect of dark matter on the shadow of a distorted and deformed compact object

### v1

@@ IV Dark matter model
The Standard Model of Cosmology characterizes the Universe as geometrically flat, with dark energy, dark matter, and ordinary matter comprising approximately $68\%$, $27\%$, and $5\%$ of its total mass-energy density, respectively. Dark matter most likely consists of unknown elementary particles beyond the standard model of particle physics that originated in the early Universe, resulting in negligible effects of its velocity dispersion on structure formation (referred to as cold dark matter). Additionally, unlike baryonic matter, dark matter is not expected to form accretion disks due to its low interaction cross-section. Furthermore, due to the absence of the coupling with the electromagnetic field, light rays may travel through dark matter distributions virtually unaffected by any non-gravitational means [^1] This is true, independently of the distribution and equation of state of the dark matter model.. To date, there has been no direct detection of dark matter. The evidence supporting its existence is primarily based on cumulative measurements, such as galactic rotation curves [60], and the cosmic microwave background [61]. In this study, we focus only on its fundamental property, more specifically its manifestation through gravitational effects [46], through the definition of an additional effective mass function as
@@
In this definition $\Delta M>0$ corresponds to the positive dark matter mass-energy density and $\Delta M<0$ to the negative one. In this work, we consider solely the positive case and replace $M$ by $m(r)$ in the line element in Eq. (II).
@@
Figures 9 show the behavior of the $g_{tt}$ and $g_{rr}$ components for the dark matter model for different combinations of $\Delta M$ and $\Delta r$. To see the effect of the dark matter distribution on the spherically symmetric black hole case, we consider the combination $\alpha=\beta=0$. Compared to the vanishing dark matter case, we see that the metric coefficients have two additional local extrema, a local minimum close to the inner edge of the disk corresponding to the inner edge of the dark matter halo, and a local maximum at the outer edge of the halo. Furthermore, the magnitude of these extrema is inversely proportional to the thickness of the halo, with large values of $\Delta r$ inducing only slight modifications to the spacetime geometry close to the central object. Furthermore, previous publications have shown that, under certain conditions, the spacetime presents an additional pair of photon spheres with opposite stability characters [46]. Note that if $r_{0}\leq r_{ph}<r_{s}$, the observer can only see the purely compact object effect on the shadow without the contribution of dark matter. Also, if $r_{s}+\Delta r<r_{ph}<r_{o}$, the dark matter halo is constrained in the vicinity of the central object but it is assumed not to fall into the event horizon, which is a rather nonphysical assumption. We thus choose to discard the two cases above and restrict our analysis to the combination $r_{s}<r_{ph}<r_{s}+\Delta r$.
@@
The methods outlined in the previous section for the ray-tracing and production of the observed intensity profiles and shadow images were also applied to the model introduced in this section. However, since the radius of the ISCO is highly sensitive to the mass distribution of the halo and it is a discontinuous function of $\Delta M$ for configurations in which the mass of the halo is much larger than the mass of the central object, we have decided to restrict our analysis to the LR and EH models for the accretion disk. The observed intensity profiles $I(r)$ for the particular case $\alpha=\beta=0$ for the LR and EH models are given in Figure 10. We chose $\alpha=\beta=0$ to analyze separately the effect of the dark matter model on the intensity profiles. Our results indicate that both the LR and the EH models are only slightly sensitive to both $\Delta M$ and $\Delta r$, with variations in the smaller values of $\Delta r$ and larger values of $\Delta M$ being the most prominent. Figure 11 shows the shadow images for the LR and EH disk models for the same particular case $\alpha=\beta=0$. This figure compares the shadow images in the absence of dark matter with two mass distributions of the same length-scale with different masses. One can observe a slight increase in the size of the shadow for both disk models with an increase in the mass of the halo.
@@
Turning now to the analysis of how the parameters $\alpha$ and $\beta$ influence the results, Figure 12 shows the observed intensity profiles for varying $\alpha$ with $\beta=0$ for a given set of parameters of dark matter model. Similarly to what was observed in the absence of the DM halo in Figures 5 and 7, variations in $\alpha$ induce radial translations of the observed intensity profiles. This is due to the fact that the accretion disk is closer to the compact object for smaller values of $\alpha$. Furthermore, upon comparing the LR panel in Figure 12 with the bottom row in Figure 5 for the cases in which $\beta=0$, it is evident that for any constant value of $\alpha$, the radial position of the peak shifts to larger values. We see the same pattern in the EH model from the comparison of Figure 12 and the bottom row in Figure 5. Furthermore, in accordance with the results of the previous section, we observe that variations in the chosen values for the parameter $\beta$ induce barely noticeable changes in the observed intensity profiles. To clarify this point, in Figure 13 we show the observed intensity profiles for $\alpha=0$ for both the DM model and in the absence of DM with varying $\beta$. These results indicate that the presence of DM does not alter the sub-dominant character of the parameter $\beta$ in the observational properties of these models. In Figure 14, we depicted the shadow images for varying $\alpha$ and vanishing $\beta$ for EH and LR models. The results are comparable with those of Figure 12.
@@
Summarizing, our results indicate that considering DM model that we chose in this work, the presence of DM halos affect only slightly the observed intensity profiles and shadows of the models considered, thus being a sub-dominant effect in comparison with the effects of the dominant deformation parameter $\alpha$. Nevertheless, for a given model with a set value of the deformation parameter, the effects of the DM halo increase with the mass and the proximity of the halo from the central compact object.
@@
Figure 11: Shadow images for the LR disk model (top row) and EH model (bottom row) with $\alpha=\beta=0$. The first column depicts the absence of dark matter, while the second and third columns show $\Delta r=50M$ and variations in $\Delta M={7M,21M}$ respectively.
@@
Figure 12: Intensity $I(r)$ as a function of the radius $r$ for the LR disk model (left plot) and for the EH disk model (right plot), with $\Delta M=20M$, $\Delta r=50M$ for $\alpha=\{-0.3,0,0.3\}$ and $\beta=0$.
Refer to caption Refer to caption
@@
Figure 14: Shadow images for the LR disk model (top row) and the EH disk model (bottom row) with $\Delta M=20M$ and $\Delta r=50M$, for $\alpha=0.3$ (left column), $\alpha=0$ (middle column), and $\alpha=-0.3$ (right column).
V Summary and conclusion
In this work, we have analyzed the shadow images and photon ring structure of a novel class of compact objects that are analytically tractable, with the goal of extending the previous research conducted in different spacetimes. This spacetime is a solution to the Einstein field equation and an extension of the familiar $\rm q$-metric, which is static and axisymmetric, and includes a distortion parameter $\beta$ and a deformation parameter $\alpha$, as briefly outlined in Section II. This analysis was done both in the absence and the presence of a dark matter halo, the latter represented by a modification in the radial mass function with two free parameters controlling the mass $\Delta M$ and the typical length-scale $\Delta r$ of the halo.
In our analysis, we have assumed the gas medium to be optically thin. The distinctive aspects of the observed images are primarily influenced by the location and characteristics of the emitting material close to the central object. In general, it can be inferred that the inclusion of quadrupoles in the metric significantly impacts the size and intensity of shadow images. Specifically, the oblateness of the central compact object, represented by the $\alpha$ parameter, plays a more pronounced role in influencing the characteristics of the shadow. This is due to the fact that modifications in $\alpha$ induce large variations in the radius of the ISCO and the photon sphere, which control the interior edge of the accretion disk models. On the other hand, we have shown that the distortion parameter $\beta$ plays a negligible role in affecting the observational properties of these spacetimes. These results indicate that the analysis of the sized of shadows observed experimentally could be used to constraint the oblateness of the central compact object, but in the absence of oblateness one can distinguish the effect of distortion due to the external matter distribution.
Furthermore, we have analyzed the effects of the presence of a dark matter halo surrounding the central object and the possibility of identifying this presence from the physical properties of emission from an optically thin disk region. Our results indicate that dark matter has the ability to either enhance or diminish the size of the observed shadow, although these effects are sub-dominant in comparison with the effects of oblateness considering the chosen parameter’s range. Note that this outcome is contingent upon the specific dark matter model being considered. In this work, we have considered a rather simple model for the dark-matter halo, however, one can consider more complicated models e.g., [62, 63, 64, 65] for further investigation.
@@ V Summary and conclusion
VI Acknowledgements
S.F. funded by the University of Waterloo. And, in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities at Perimeter Institute. This work is initiated at ZARM, Germany. J.L.R. acknowledges the European Regional Development Fund and the programme Mobilitas Pluss for financial support through Project No. MOBJD647, project No. 2021/43/P/ST2/02141 co-funded by the Polish National Science Centre and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Sklodowska-Curie grant agreement No. 94533, Fundação para a Ciência e Tecnologia through project number PTDC/FIS-AST/7002/2020, and Ministerio de Ciencia, Innovación y Universidades (Spain), through grant No. PID2022-138607NB-I00.

### v2

@@ 4 Dark matter model
The Standard Model of Cosmology characterizes the Universe as geometrically flat, with dark energy, dark matter, and ordinary matter comprising approximately $68\%$, $27\%$, and $5\%$ of its total mass-energy density, respectively. Dark matter most likely consists of unknown elementary particles beyond the standard model of particle physics that originated in the early Universe, resulting in negligible effects of its velocity dispersion on structure formation (referred to as cold dark matter). Additionally, unlike baryonic matter, dark matter is not expected to form accretion disks due to its low interaction cross-section. Furthermore, due to the absence of the coupling with the electromagnetic field, light rays may travel through dark matter distributions virtually unaffected by any non-gravitational means [^3] This is true, independently of the distribution and equation of state of the dark matter model.. To date, there has been no direct detection of dark matter. The evidence supporting its existence is primarily based on cumulative measurements, such as galactic rotation curves (Rubin et al., 1980), and the cosmic microwave background (Planck Collaboration, 2014). In this study, we focus only on its fundamental property, more specifically, its manifestation through gravitational effects (Konoplya, 2019), through the definition of an additional effective mass function as
@@
In this definition, $\Delta M>0$ corresponds to the positive dark matter mass-energy density and $\Delta M<0$ to the negative one. In this work, we consider solely the positive case and replace $M$ by $m(r)$ in the line element in Eq. (2).
@@
Figure 6 shows the behaviour of the $g_{tt}$ and $g_{rr}$ components for the dark matter model for different combinations of $\Delta M$ and $\Delta r$. To see the effect of the dark matter distribution on the spherically symmetric black hole case, we consider the combination $\alpha=\beta=0$. Compared to the vanishing dark matter case, we see that the metric coefficients have two additional local extrema, a local minimum close to the inner edge of the disk corresponding to the inner edge of the dark matter halo, and a local maximum at the outer edge of the halo. Furthermore, the magnitude of these extrema is anti-correlated to the thickness of the halo, with large values of $\Delta r$ inducing only slight modifications to the spacetime geometry close to the central object. Furthermore, previous publications have shown that, under certain conditions, the spacetime presents an additional pair of unstable photon orbits with opposite stability characters (Konoplya, 2019). Note that if $r_{o}\leq r_{LR}<r_{s}$, where $r_{o}$ is the radius at which the observer stands, the observer can only see the purely compact object effect on the shadow without the contribution of dark matter. Also, if $r_{s}+\Delta r<r_{LR}<r_{o}$, the dark matter halo is constrained in the vicinity of the central object but it is assumed not to fall into the event horizon, which is a rather nonphysical assumption. We thus choose to discard the two cases above and restrict our analysis to the combination $r_{s}<r_{LR}<r_{s}+\Delta r$.
@@
The methods outlined in the previous section for the ray-tracing and production of the observed intensity profiles and shadow images were also applied to the model introduced in this section. However, since the radius of the ISCO is highly sensitive to the mass distribution of the halo and it is a discontinuous function of $\Delta M$ for configurations in which the mass of the halo is much larger than the mass of the central object, we have decided to restrict our analysis to the LR and EH models for the accretion disk. The observed intensity profiles $I(r)$ for the particular case $\alpha=\beta=0$ for the LR and EH models are given in Figure 7. We chose $\alpha=\beta=0$ to analyse separately the effect of the dark matter model on the intensity profiles. Our results indicate that both the LR and the EH models are only slightly sensitive to both $\Delta M$ and $\Delta r$, with variations in the smaller values of $\Delta r$ and larger values of $\Delta M$ being the most prominent. Figure 8 shows the shadow images for the LR and EH disk models for the same particular case $\alpha=\beta=0$. This figure compares the shadow images in the absence of dark matter with two mass distributions of the same length-scale with different masses. One can observe a slight increase in the size of the shadow for both disk models with an increase in the mass of the halo.
@@
Turning now to the analysis of how the parameters $\alpha$ and $\beta$ influence the results, Figure 9 shows the observed intensity profiles for varying $\alpha$ with $\beta=0$ for a given set of parameters of the dark matter model. Similarly to what was observed in the absence of the DM halo in Figure 3 and Figure 4, variations in $\alpha$ induce radial translations of the observed intensity profiles. This is due to the fact that the accretion disk is closer to the compact object for smaller values of $\alpha$. Furthermore, upon comparing the LR panel in Figure 4 and Figure 9 for the cases in which $\beta=0$, it is evident that for any constant value of $\alpha$, the radial position of the peak shifts to larger values. We see the same pattern in the EH model panels of Figure 9 and Figure 4. Furthermore, in accordance with the results of the previous section, we observe that variations in the chosen values for the parameter $\beta$ induce barely noticeable changes in the observed intensity profiles. To clarify this point, in Figure 10 we show the observed intensity profiles for $\alpha=0$ for both the DM model and in the absence of DM with varying $\beta$. These results indicate that the presence of DM does not alter the sub-dominant character of the parameter $\beta$ in the observational properties of these models. In Figure 11, we depicted the shadow images for varying $\alpha$ and vanishing $\beta$ for EH and LR models. The results are comparable with those of Figure 9.
@@
Summarising, our results indicate that considering DM model that we chose in this work, the presence of DM halos affects only slightly the observed intensity profiles and shadows of the models considered, thus being a sub-dominant effect in comparison with the effects of the dominant deformation parameter $\alpha$. Nevertheless, for a given model with a set value of the deformation parameter, the effects of the DM halo increase with the mass and the proximity of the halo from the central compact object.
@@
Figure 8: Shadow images for the LR disk model (top row) and EH model (bottom row) with $\alpha=\beta=0$. The first column depicts the absence of dark matter, while the second and third columns show $\Delta r=50M$ and variations in $\Delta M={7M,21M}$, respectively.
@@
Figure 9: Intensity $I(r)$ as a function of the radius $r$ for the LR disk model (left plot) and for the EH disk model (right plot), with $\Delta M=20M$, $\Delta r=50M$ for $\alpha=\{-0.01,0,0.01\}$ and $\beta=0$.
@@
Figure 11: Shadow images for the LR disk model (top row) and the EH disk model (bottom row) with $\Delta M=20M$ and $\Delta r=50M$, for $\alpha=0.01$ (left column), $\alpha=0$ (middle column), and $\alpha=-0.01$ (right column).
5 Summary and Conclusion
In this work, we have analysed the shadow images and photon ring structure of a novel class of compact objects that are analytically tractable, with the goal of extending the previous research conducted in different spacetimes. This spacetime is a solution to the Einstein field equation and an extension of the familiar $\rm q$-metric, which is static and axisymmetric, and includes a distortion parameter $\beta$ and a deformation parameter $\alpha$, as briefly outlined in Section 2. This analysis was done both in the absence and the presence of a dark matter halo, the latter represented by a modification in the radial mass function with two free parameters controlling the mass $\Delta M$ and the typical length-scale $\Delta r$ of the halo.
In our analysis, we have assumed the gas medium to be optically thin. The distinctive aspects of the observed images are primarily influenced by the location and characteristics of the emitting material close to the central object. In general, it can be inferred that the inclusion of a quadrupole in the metric significantly impacts the size and intensity of shadow images. Specifically, the oblateness of the central compact object, represented by the $\alpha$ parameter, plays a more pronounced role in influencing the characteristics of the shadow. This is due to the fact that modifications in $\alpha$ induce large variations in the radius of the ISCO and the unstable photon orbit, which control the interior edge of the accretion disk models. On the other hand, we have shown that the distortion parameter $\beta$ plays a negligible role in affecting the observational properties of these spacetimes. These results indicate that the analysis of the sizes of shadows observed experimentally could be used to constraint the oblateness of the central compact object, but in the absence of oblateness, one can distinguish the effect of distortion due to the external matter distribution.
Furthermore, we have analysed the effects of the presence of a dark matter halo surrounding the central object and the possibility of identifying this presence from the physical properties of emission from an optically thin disk region. Our results indicate that dark matter has the ability to either enhance or diminish the size of the observed shadow, although these effects are sub-dominant in comparison with the effects of oblateness, considering the chosen parameter’s range. Note that this outcome is contingent upon the specific dark matter model being considered. In this work, we have considered a rather simple model for the dark-matter halo, however, one can consider more complicated models e.g., (Hernquist, 1990; Kiselev, 2003; Li & Yang, 2012; Xu et al., 2020) for further investigation.
@@ 5 Summary and Conclusion
Acknowledgements.
Sh.F. is funded by the University of Waterloo and in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities at Perimeter Institute. Also, thanks the excellence cluster QuantumFrontiers funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers. J.L.R. acknowledges the European Regional Development Fund and the programme Mobilitas Pluss for financial support through Project No. MOBJD647, project No. 2021/43/P/ST2/02141 co-funded by the Polish National Science Centre and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Sklodowska-Curie grant agreement No. 94533, Fundação para a Ciência e Tecnologia through project number PTDC/FIS-AST/7002/2020, and Ministerio de Ciencia, Innovación y Universidades (Spain), through grant No. PID2022-138607NB-I00.
Appendix A Validity of the spherical symmetry approximation
In this section, we prove that the approximation taken in this work, i.e., that one can integrate the geodesic equation and obtain the trajectory of the light rays by taking the spacetime to be approximately spherically symmetric. For this purpose, and since the geodesic equation for the photons depends on the $g_{rr}$, $g_{tt}$, and $g_{\phi\phi}$ components of the metric, we calculate the relative errors between these metric components of the spherically symmetric Schwarzschild solution (defined by the parameters $\alpha=0$ and $\beta=0$) and those of the solutions with non-zero $\alpha$ and $\beta$. We define the relative error as:
$\nu_{i}=\left|\frac{g_{ii}^{S}-g_{ii}}{g_{ii}^{S}}\right|,$ (17)
where $\nu$ is the relative error, the subscript $i$ can denote either $r$, $t$, or $\phi$, and $g_{\mu\nu}^{S}$ is the Schwazrschild spacetime metric. In the region of interest where photons are propagating between the observer and the central object and not falling into the event horizon, i.e., for a range $r\in\left[3M,100M\right]$ and $\theta=\left[0,\frac{\pi}{2}\right]$, the relative errors are plotted in Figure 12. These results show that, in the regions where the photons propagate, the relative errors are consistently smaller than $1\%$, thus validating the approximation used.
Figure 12: Relative errors $\nu_{r}$ (top row), $\nu_{t}$ (middle row), and $\nu_{\phi}$ (bottom row) as defined in Eq. (17) for $\alpha=\mp 0.01$ and $\beta=\pm 10^{-6}$ (left column), $\alpha=0$ and $\beta=\pm 10^{-6}$, and $\alpha=\pm 0.01$ and $\beta=\pm 10^{-6}$. The shades represent, from darker to lighter, the regions where the errors are below $0.02\%$, $0.05\%$, $0.1\%$, $0.2\%$, $0.5\%$, and $1\%$, with the largest errors being of the order of $2\%$.
Appendix B Complete shadow images and intensity plots
For completeness, in this appendix we show the complete set of shadow images and intensity plots for all of the parameter combinations, i.e., $\alpha=\{-0.01,0,0.01\}$ and $\beta=\{-10^{-5},0,10^{-5}\}$, and for all of the accretion disk models, i.e., the ISCO, the LR, and the EH model in Figures 13-18. These results serve to illustrate the negligible effect of the parameter $\beta$ to the observable properties of the system.
Figure 13: Shadow images for the ISCO disk model with different combinations of $\alpha$ and $\beta$, namely $\alpha=-0.01$ (top row), $\alpha=0$ (middle row), and $\alpha=+0.01$ (bottom row); and $\beta=-10^{-5}$ (left column), $\beta=0$ (middle column), and $\beta=10^{-5}$ (right) column).
Figure 15: Shadow images for the LR disk model with different combinations of $\alpha$ and $\beta$, namely $\alpha=-0.01$ (top row), $\alpha=0$ (middle row), and $\alpha=+0.01$ (bottom row); and $\beta=-10^{-5}$ (left column), $\beta=0$ (middle column), and $\beta=10^{-5}$ (right) column).
Figure 17: Shadow images for the EH disk model with different combinations of $\alpha$ and $\beta$, namely $\alpha=-0.01$ (top row), $\alpha=0$ (middle row), and $\alpha=+0.01$ (bottom row); and $\beta=-10^{-5}$ (left column), $\beta=0$ (middle column), and $\beta=10^{-5}$ (right) column).
Figure 18: Intensity $I(r)$ as a function of the radius $r$ for the EH disk model for a varying $\alpha$ with constant $\beta$ (top row) and for a varying $\beta$ with constant $\alpha$ (bottom row).